OpenStudy (anonymous):

line that passes through the point (4, -6) and is parallel to y = -1 over 2 x + 5

OpenStudy (jh3power):

well since it's parallel, it should be y=a kind of graph. so it's y=-6.

OpenStudy (jh3power):

what was 'over 2x+5' for?

OpenStudy (anonymous):

i put -1 over 2 then it was x+5

OpenStudy (anonymous):

\[ y = \frac{-1}{2}( x + 5)\] comparing above eq with y=mx+c we find slope m= -1/2 since line passes through point (4, -6) therefore eq of line in slope point form is given as y-y1=m(x-x1) i.e. y+6=-1/2(x-4) i.e. y+6=-1/2x+2 i.e. y=-1/2x+2 -6 i.e. y=-1/2x-4 2y =-x-8 i.e. x+2y =-8 which is the required eq of line

OpenStudy (jh3power):

oh.. i thought it was just y=-1

OpenStudy (anonymous):

y = -2x + 2 y = 2x - 14 dese are the answers and the rest of the asnwers are y= 1 over 2